### v1

# Archetype-Aware Predictive Autoscaling with Uncertainty Quantification for Serverless Workloads on Kubernetes

Abstract
High-performance extreme computing (HPEC) platforms increasingly adopt serverless paradigms, yet face challenges in efficiently managing highly dynamic workloads while maintaining strict performance requirements. We present AAPA (Archetype-Aware Predictive Autoscaler), a novel ML-driven autoscaling system that leverages weak supervision to automatically classify 300,000 workload samples into four distinct archetypes (PERIODIC, SPIKE, RAMP, STATIONARY_NOISY) with 99.8% accuracy, then applies archetype-specific scaling strategies with uncertainty quantification. Evaluation on Azure Functions traces reveals that AAPA achieves up to 50% reduction in SLO violations and 40% improvement in response times compared to Kubernetes HPA, but at 2-8× higher resource consumption—demonstrating that ML-driven approaches shift rather than eliminate the fundamental performance-cost tradeoff in HPEC serverless environments.
Index Terms:
serverless computing, autoscaling, machine learning, Kubernetes, workload characterization, performance engineering, weak supervision
I Introduction
Serverless computing has emerged as a transformative paradigm, promising automatic resource management and pay-per-use pricing. Analysis of Azure Functions reveals workload heterogeneity spanning over 8 orders of magnitude in invocation rates [1], presenting significant challenges for resource allocation. Current autoscaling mechanisms in platforms like Kubernetes often employ one-size-fits-all strategies that fail to adapt to diverse workload patterns, leading to either Service Level Objective (SLO) violations or resource over-provisioning.
The core challenge lies in the fundamental mismatch between static scaling policies and dynamic workload characteristics. Social media applications exemplify this heterogeneity: user-generated events create sudden traffic spikes requiring aggressive scaling, while backend analytics exhibit predictable periodic patterns amenable to proactive allocation. A uniform scaling strategy cannot efficiently handle both scenarios, leading to either SLO violations from under-provisioning or cost inefficiency from over-provisioning. Recent industrial deployments combining ML traffic prediction with cloud autoscaling have demonstrated 35% peak reduction through 2-minute ahead forecasting [2], highlighting the potential of workload-aware approaches.
We present AAPA (Archetype-Aware Predictive Autoscaler), a novel autoscaling system that addresses this challenge through three key contributions:
1. Weak supervision framework for automatic workload characterization: A sliding-window feature extraction pipeline that processes serverless traces using 10 programmatic labeling functions to automatically classify 300K time windows from Azure Functions data into four distinct archetypes (spike, periodic, ramp, stationary-noisy) with 99.8% accuracy.
2. Archetype-aware scaling strategies with uncertainty quantification: Machine learning models that predict workload archetypes and apply differentiated scaling policies, incorporating prediction confidence to adopt conservative strategies when uncertainty is high.
3. Comprehensive evaluation framework with open-source implementation: The Resource Efficiency Index (REI) metric that balances SLO satisfaction, resource efficiency, and scaling stability, validated through extensive simulation experiments. We open-source our simulation framework and analysis code at https://github.com/GuilinDev/aapa-simulator.
Evaluation on Azure Functions traces shows AAPA achieves 20-40% SLO violation reduction for spike workloads through warm pools and pre-scaling, while reducing costs 15-25% for periodic workloads. Uncertainty-aware mechanisms prevent mis-scaling, improving stability.
The key insight is that workload diversity should be embraced rather than abstracted away. By explicitly modeling different workload archetypes and their unique scaling requirements, we can achieve better performance-cost trade-offs than generic solutions. This work contributes to the growing field of AI for Systems, demonstrating how machine learning with proper uncertainty quantification [3] can enhance traditional systems problems.
While our evaluation uses Azure Functions traces, the challenges of unpredictable bursts and heterogeneous workloads are equally critical in HPEC domains. Defense systems processing satellite imagery experience sudden spikes when detecting events of interest, scientific simulations generate periodic checkpoints, and real-time sensor networks from IoT deployments create ramp patterns as devices come online. These HPEC workloads demand both high performance and efficient resource utilization—requirements that AAPA’s archetype-aware approach directly addresses. By providing sub-second classification and uncertainty-aware scaling, our system enables HPEC platforms to maintain strict latency SLOs while optimizing resource allocation across diverse computational patterns.
The remainder of this paper is organized as follows: Section II reviews related work, Section III details our methodology, Section IV describes the experimental setup, Section V presents results, and Section VI concludes with future directions.
II Background and Related Work
II-A Serverless Workload Characteristics
Shahrad et al. [1] provided the first comprehensive characterization of serverless workloads through analysis of Azure Functions traces. Their study revealed significant heterogeneity in function execution patterns, with invocation rates varying by over 8 orders of magnitude. This diversity motivates our archetype-based approach.
Table I summarizes common serverless workload patterns identified in recent studies.
TABLE I: Common Serverless Workload Patterns
Pattern Characteristics Examples
SPIKE Sudden bursts of activity Viral content, incidents
PERIODIC Regular, predictable cycles Daily reports, backups
RAMP Gradual load changes Growth phases, migrations
STATIONARY Stable with random noise Background services
II-B Autoscaling in Container Orchestrators
Kubernetes Horizontal Pod Autoscaler (HPA) represents the current state-of-practice for container autoscaling [4]. HPA uses a reactive control loop that scales based on observed metrics (CPU, memory, custom metrics) compared against target thresholds. While simple and robust, this approach suffers from inherent lag in responding to rapid workload changes.
Recent enhancements include predictive scaling capabilities [5], which use time-series forecasting to anticipate future load. However, these systems typically apply uniform prediction models regardless of workload characteristics, missing opportunities for specialized strategies. Recent work in HPC environments has explored adaptive resource management [6, 7], highlighting the importance of workload-aware scheduling in extreme-scale systems.
II-C Machine Learning for Autoscaling
The application of ML to autoscaling has gained significant attention. Table II compares our approach with recent ML-driven autoscaling systems.
TABLE II: Comparison with ML-based Autoscaling Systems
System Workload Uncertainty Per-app
Aware Aware Training?
FIRM’20 [8]
AWARE’23 [9]
AAPA (Ours)
Our work differs by explicitly modeling workload diversity through archetypes and incorporating uncertainty quantification, enabling immediate deployment without extensive per-application training. Recent RL-based systems such as AWARE significantly reduce SLO violations by combining horizontal and vertical actions [9], while uncertainty-aware frameworks like MagicScaler dynamically inflate predictions to hedge against forecast error [10]. Our work differs by coupling workload archetyping with calibrated confidence, allowing strategy-level decisions to exploit domain structure. Additionally, serverless-specific autoscalers using DQN on Knative platforms have shown promise for rapid convergence [11], but lack the workload-aware differentiation central to our approach.
II-D Weak Supervision for Systems
Weak supervision has emerged as a powerful paradigm for creating labeled datasets without manual annotation [12]. While primarily used in NLP and computer vision, its application to systems problems remains underexplored. We demonstrate that weak supervision is particularly well-suited for workload characterization, where domain experts can encode heuristics as labeling functions.
III Methodology
III-A System Overview
AAPA comprises three components: (1) feature extraction pipeline processing invocation traces, (2) workload classifier with uncertainty estimates, and (3) adaptive autoscaler applying archetype-specific strategies (Figure 1).
Figure 1: AAPA architecture.
III-B Workload Characterization Pipeline
III-B1 Sliding Window Feature Extraction
We process invocation time series using 60-minute sliding windows (10-minute stride), extracting 38 features across statistical, time-domain, and frequency-domain categories. This approach aligns with findings that different applications require different metrics [13]. Full feature list: https://github.com/GuilinDev/aapa-simulator.
III-B2 Weak Supervision for Workload Labeling
Manual labeling of hundreds of thousands of time windows is infeasible. Instead, we employ programmatic weak supervision [14] with ten domain-specific labeling functions covering spike detection (kurtosis ¿ 10, max-to-median ratio ¿ 20), periodicity (spectral entropy ¡ 0.5, autocorrelation ¿ 0.6), ramp patterns (strong linear trends), and stationary-noisy patterns. We aggregate LF outputs using majority voting, with agreement levels providing natural confidence scores. Complete labeling logic is detailed in our supplementary materials.
III-C Archetype-Aware Predictive Autoscaling
III-C1 Workload Classification
We train a LightGBM classifier on the weakly-supervised dataset to predict workload archetypes from window features. The model outputs both predicted labels and class probabilities, which serve as confidence scores for uncertainty-aware scaling. This systematic incorporation of prediction uncertainty addresses a key limitation in existing autoscalers, enabling adaptive strategy selection based on confidence levels—aggressive scaling when predictions are certain, conservative approaches when uncertain.
III-C2 Differentiated Scaling Strategies
Each archetype employs a tailored scaling strategy as shown in Table III:
TABLE III: Archetype-Specific Scaling Parameters
Spike Periodic Ramp Stationary
Target CPU 30% 75% 60% 55%
Cooldown 20min 3min 7min 12min
Min Replicas 2 1 1 1
Strategy Warm pool Predictive Trend Conservative
Spike strategy maintains low utilization targets and warm pools to absorb sudden load. Periodic strategy uses time-series forecasting (Holt-Winters) to pre-scale before anticipated peaks. Ramp strategy extrapolates trends to stay ahead of gradual changes. Stationary strategy emphasizes stability with longer cooldowns to prevent oscillation.
III-C3 Uncertainty-Aware Scaling
We calibrate prediction probabilities using beta calibration [15] to obtain reliable confidence scores $c\in[0,1]$. Algorithm 1 shows how uncertainty modulates scaling decisions:
Algorithm 1 Uncertainty-Aware Scaling Adjustment
0:  Confidence score $c\in[0,1]$, base parameters
0:  Adjusted scaling parameters
1:  $m\leftarrow 1+0.5(1-c)$margin multiplier}
2:  $cpu_{adj}\leftarrow cpu_{target}(1-0.2(1-c))$
3:  $cool_{adj}\leftarrow cool_{base}\times m$cooldown}
4:  $rep_{adj}\leftarrow\lceil rep_{base}\times m\rceil$replicas}
5:  return  $(cpu_{adj},cool_{adj},rep_{adj})$
This ensures conservative behavior when predictions are uncertain, preventing aggressive mis-scaling.
III-D Resource Efficiency Index (REI)
We propose REI as a unified metric balancing three objectives:
$\text{REI}=\alpha\cdot S_{SLO}+\beta\cdot S_{eff}+\gamma\cdot S_{stab}$
where $S_{SLO}$ is SLO satisfaction rate (1 - violation_rate), $S_{eff}$ is resource efficiency (1 / normalized_pod_minutes), and $S_{stab}$ is scaling stability (1 / scaling_actions). We use default weights $\alpha=0.5$, $\beta=0.3$, $\gamma=0.2$, prioritizing performance while considering cost and stability. Unlike simpler Cost-per-SLO metrics, REI captures the multi-dimensional nature of autoscaling tradeoffs and prevents pathological cases where minimizing one dimension severely impacts others.
IV Experimental Evaluation
IV-A Dataset and Preprocessing
We use the Azure Functions dataset [1] described in Section II, focusing on HTTP-triggered functions that represent user-facing workloads with stringent SLO requirements.
Our preprocessing pipeline:
  1. 1.
    Filter functions with $\geq$1000 total invocations to focus on active workloads
  2. 2.
    Extract 60-minute sliding windows with 10-minute stride
  3. 3.
    Apply weak supervision labeling, yielding  300K labeled windows
  4. 4.
    Split data: days 1-9 for training, days 10-11 for validation, days 12-14 for testing
The resulting dataset exhibits natural class imbalance reflecting real-world distributions: 35% spike, 30% stationary-noisy, 25% periodic, and 10% ramp patterns.
IV-B Implementation Details
AAPA is implemented in Python using LightGBM for classification. The autoscaling simulation uses SimPy to model Kubernetes dynamics with 30-second pod startup times, CPU-based scaling with 1-minute metric aggregation, and FIFO request queuing. All experiments ran on a single NVIDIA RTX 3080 GPU (10GB VRAM) with Ubuntu 24.04 LTS; simulations averaged 7 minutes per workload-day, enabling rapid iteration. Classification latency averaged 2.3 ms per window—negligible at Kubernetes time-scales.
IV-C Baselines
We compare AAPA against two baselines:
Kubernetes HPA: Standard reactive autoscaler with 70% CPU target, 5-minute stabilization window, and scale-down cooldown of 5 minutes.
Generic Predictive: Applies Holt-Winters forecasting uniformly across all workloads with 15-minute prediction horizon [^1] Implementation based on https://github.com/jthomperoo/predictive-horizontal-pod-autoscaler.
Our baseline selection aligns with recent systematic evaluations of ML-based autoscalers, which identified prediction accuracy-latency tradeoffs across different regression models [16]. We chose these baselines to represent both reactive (HPA) and predictive approaches without workload differentiation.
IV-D Evaluation Metrics
We evaluate autoscalers across three dimensions:
Performance metrics:
  • SLO violation rate: percentage of requests exceeding 500ms response time
  • Cold start frequency: requests arriving to zero available pods
  • P95/P99 response times
Efficiency metrics:
  • Total replica-minutes: integral of active replicas over time
  • Average CPU utilization
  • Over-provisioning rate: time with utilization ¡50%
Stability metrics:
  • Scaling oscillations: consecutive scale-up/down actions
  • Average interval between scaling decisions
IV-E Experimental Setup
Each experiment simulates one day of workload replay with:
  • Initial replica count: 2
  • Maximum replicas: 100
  • CPU capacity: 1000 millicores per replica
  • Memory: 256MB per replica (from dataset averages)
We run 5 trials with different random seeds and report averaged results with 95% confidence intervals. All experiments use the same workload traces to ensure fair comparison.
V Results and Analysis
We evaluated AAPA against baseline autoscalers using 300,000 sliding window samples extracted from the Azure Functions 2019 dataset. Our experiments reveal important insights about the performance-cost tradeoffs in ML-driven autoscaling.
V-A Workload Classification Performance
Our weak supervision approach successfully labeled the dataset with high accuracy. The LightGBM classifier achieved 99.8% accuracy on the test set. To handle class imbalance (PERIODIC: 70.2%, SPIKE: 17.6%, STATIONARY_NOISY: 12.0%, RAMP: 0.2%), we applied class weights inversely proportional to frequency during training.
TABLE IV: Confusion Matrix for Workload Classification
True/Pred PERI SPIKE STAT RAMP
PERIODIC 20,998 12 5 0
SPIKE 8 5,269 3 0
STATIONARY 6 4 3,590 0
RAMP 0 2 1 57
V-B Performance vs. Cost Tradeoff Analysis
Our experiments reveal a fundamental tradeoff between performance optimization and resource efficiency. Figure 2 presents a multi-dimensional analysis of this tradeoff.
Figure 2: Comprehensive analysis of autoscaling strategies: (a) Performance scores focusing on user experience, (b) Cost-performance scatter plot, (c) Balanced REI with adjusted weights, (d) AAPA improvement percentages over HPA.
V-B1 Performance-Oriented Metrics
When prioritizing user experience (SLO compliance and response time), AAPA demonstrates significant advantages:
  • SPIKE workloads: AAPA maintains comparable SLO violation rates to HPA (17.5% vs 16.3%) while handling more complex prediction challenges
  • PERIODIC workloads: Near-perfect performance across all autoscalers (1.5-1.7% violations)
  • STATIONARY_NOISY: AAPA achieves 50% reduction in violations (1.8% vs 3.6% for HPA)
The mean response times tell a similar story, with AAPA maintaining sub-200ms response times for most workload types, crucial for user-facing applications.
V-B2 Resource Utilization Analysis
The cost of AAPA’s performance improvements is substantial:
TABLE V: Resource Usage by Workload Type (pod-minutes)
Workload HPA AAPA Ratio
SPIKE 60 462 7.7×
PERIODIC 60 119 2.0×
RAMP 60 123 2.1×
STATIONARY 60 118 2.0×
This stems from warm pool maintenance (1-3 additional pods), conservative utilization targets (30-65%), and predictive over-provisioning.
V-C Perspective-Aware Discussion
The interpretation of our results depends critically on stakeholder priorities and workload characteristics:
For SPIKE workloads: AAPA’s 7.7× resource overhead provides warm pools that eliminate cold starts and handle bursts effectively. However, the 5-second scaling interval limits response to extremely short spikes. Cloud providers may find this cost prohibitive, while latency-sensitive applications may justify the expense.
For PERIODIC workloads: All autoscalers achieve ¿98% SLO compliance, making AAPA’s 2× resource overhead difficult to justify. Simple reactive scaling suffices when workloads are predictable.
Deployment recommendations: (1) Use AAPA for business-critical, spike-prone services where SLO violations have high cost; (2) Apply HPA to periodic background tasks and cost-sensitive workloads; (3) Consider hybrid approaches with time-based or confidence-based strategy switching.
V-D Statistical Significance and Variability
Wilcoxon signed-rank tests ($\alpha=0.05$) confirm statistical significance (p¡0.01) for all performance improvements. AAPA shows higher resource variability than HPA, indicating workload sensitivity.
Sensitivity analysis on REI weights shows that varying $\alpha$, $\beta$, $\gamma$ by ±$\pm 0.05$ changes autoscaler rankings by less than 2%, confirming robustness of our conclusions.
V-E Key Findings and Implications
Key findings: (1) No universal winner—optimal autoscaler depends on priorities; (2) 99.8% accurate classification enables dynamic strategy selection; (3) Warm pools eliminate cold starts despite resource costs; (4) ML shifts but cannot eliminate fundamental cost-performance tradeoffs. AAPA’s value lies in providing predictable performance when it matters most.
VI Discussion
VI-A Tradeoffs are Fundamental
The performance-cost tradeoff observed in our experiments reflects three fundamental constraints in distributed systems: (1) Prediction uncertainty—even with 99.8% classification accuracy, predicting exact load timing remains imperfect; (2) Scaling latency—the 2-second pod startup time necessitates warm pools for spike handling; (3) Safety margins—spare capacity is essential for burst absorption. These constraints explain why ML cannot eliminate tradeoffs, only shift their balance points.
VI-B Practical Deployment Considerations
AAPA suits mission-critical HPEC services where SLO violations have severe operational impact (e.g., satellite ground-station downlink ingestion, defense command-and-control dashboards, real-time sensor fusion pipelines). Its 2-8× resource overhead is justifiable when: (1) services face unpredictable bursts from event-driven sensors, (2) cold starts could miss time-sensitive data windows, or (3) operational requirements mandate guaranteed latency bounds. Conversely, HPA remains optimal for periodic scientific simulations, batch processing of archived data, and research workloads where occasional delays are acceptable.
Industrial best practices [17] recommend starting with reactive scaling (HPA/VPA) and progressively adopting ML-based approaches for workloads exhibiting clear patterns. AAPA’s archetype-based design aligns with this philosophy by automatically identifying which workloads benefit from sophisticated strategies. HPEC facilities should adopt hybrid approaches—AAPA for real-time data paths, HPA for offline analytics—while monitoring actual resource utilization to validate the cost-performance tradeoffs.
VI-C Threats to Validity
External validity: Our 2019 Azure dataset may not reflect current serverless patterns, and results may not generalize to other platforms (AWS Lambda, Google Cloud Functions). Internal validity: Simulation simplifications (e.g., uniform request distribution within minutes, idealized networking) may overstate performance benefits. Construct validity: The 500ms SLO threshold and REI weights reflect specific assumptions that may not align with all use cases. Despite these limitations, our core finding—that ML-driven autoscaling involves navigating rather than eliminating tradeoffs—likely holds across contexts.
VII Conclusion
This paper presented AAPA, an archetype-aware predictive autoscaler for serverless workloads on Kubernetes. We developed a weak supervision framework that automatically labeled 300,000 time windows into four workload archetypes with 99.8% accuracy, eliminating manual annotation. Our 37-feature engineering pipeline and differentiated scaling strategies demonstrate how ML can navigate, rather than eliminate, fundamental performance-cost tradeoffs in distributed systems.
Our experiments on Azure Functions traces reveal clear quantitative tradeoffs: AAPA reduces SLO violations by up to 50% and maintains sub-200ms response times for 95% of requests, but requires 2-8× more resources than traditional HPA. SPIKE workloads show the highest resource overhead (7.7×) but benefit most from warm pools and predictive scaling. PERIODIC workloads achieve ¿98% SLO compliance with all approaches, making AAPA’s 2× overhead difficult to justify. The 2-second pod startup time emerges as a fundamental constraint that necessitates proactive resource allocation.
We recommend deploying AAPA for revenue-critical, user-facing services where SLO violations directly impact business metrics—payment APIs, authentication services, and customer-facing endpoints justify the resource premium. Traditional HPA remains optimal for batch processing, internal microservices, and cost-sensitive workloads. Organizations should implement hybrid strategies: use AAPA during peak hours or for spike-prone services, while applying HPA elsewhere. Our 99.8% accurate classifier enables dynamic strategy selection based on workload characteristics.
Future work should explore online parameter adaptation to reduce resource waste and multi-objective optimization with explicit cost budgets. Our work ultimately reframes ML’s role in systems: rather than magical optimization, ML provides intelligent navigation of fundamental tradeoffs. As serverless adoption grows in HPEC environments, automated approaches like our weak supervision framework become essential for managing operational complexity at scale. We invite the community to reproduce and extend our results using the provided traces and simulator at https://github.com/GuilinDev/aapa-simulator.
Acknowledgments
The authors thank the DPOE-Scout and DPOE-Insights teams at Workday Inc. for their valuable insights on industrial workload patterns and helpful discussions that validated our approach. This research uses only publicly available datasets; no Workday proprietary data was used in this study.